OpenStudy (anonymous):

Solve the equation for y: 4x + 3y = -12.

OpenStudy (anonymous):

4x + 3y = -12 4x = -12 - 3y So, x = -12-3y/4 Replace x in the equation and you will get the answer.

OpenStudy (king):

take x=0 then y=-4